you row a boat upsteam at 4 mph, down steam at 6 mph and it takes you 3 hours total, how many miles did you travel in 3 hours
Answer by richwmiller(17219)

stream not steam
x+y=3
4x=6y
x=9/5 y=6/5
4*9/5+6*6/5=d
36/5+36/5=72/5
14 2/5 miles